OpenStudy (anonymous):

The graph of the following system of equations is 2x + y = 6 6x + 3y = 18 i think its overlapping....

OpenStudy (anonymous):

@superhelp101 or parallel can you help? these are the options overlapping lines parallel lines intersecting lines

OpenStudy (superhelp101):

yeah i would choose overlapping too but i was thinking parallel at first, u could put this in your fancy graphing calculator to make sure, i would for u but my calculator ran out of batteries :''(

OpenStudy (anonymous):

do you know a good site for that @superhelp101

OpenStudy (superhelp101):

good site for what?

OpenStudy (anonymous):

graphing

OpenStudy (superhelp101):

oh yea i think that there is a cite but i remember i was trying to find it soo hard but i couldn't sooo sorry

OpenStudy (anonymous):

its ok:) can you help with a couple more @superhelp101

OpenStudy (superhelp101):

sure go ahead

OpenStudy (anonymous):

Which of the following statements best describes the graph of x + y = 3? It is a line joining the points whose x and y-coordinates add up to 3. It is a line joining the points whose x and y-coordinates add up to 6. It is a line which intersects the x-axis at (3, 3). It is a line which intersects the y-axis at (3, 3).

OpenStudy (superhelp101):

u gotta know the Slope-intercept formula y=mx+b 2. y = -x + 5 answer is It is a line joining the points whose x and y coordinates add up to 5.

OpenStudy (anonymous):

A?

OpenStudy (anonymous):

thats what i thought it was

OpenStudy (superhelp101):

yes, good job!

OpenStudy (anonymous):

am i correct @superhelp101

OpenStudy (superhelp101):

alright for this one i am not to sure but i highly think so...

OpenStudy (anonymous):

@superhelp101 ok :) can you continue to assist please you don't have too

OpenStudy (superhelp101):

i'll try (:

OpenStudy (anonymous):

A company makes rugs in two sizes. A small size rug requires 2 hours for dyeing and 4 hours for knotting. A medium size rug requires 3 hours for dyeing and 6 hours for knotting. The company's dyeing unit cannot work for more than 40 hours in a week, and the knotting unit cannot work for more than 50 hours in week. Let S represent the number of small size rugs that the company can manufacture in a week and M represent the number of medium size rugs that the company can manufacture in a week. Which of the following statements best explains whether the company will be able to make 12 small size rugs and 5 medium size rugs in a week? The company will not be able to make these numbers in a week because of the system restrictions of 2S + 3M ≤ 40 and 4S + 6M ≤ 50. The company will be able to make these numbers in a week because of the system restrictions of 2S + 4S ≤ 40 and 3M + 6M ≤ 50. The company will be able to make these numbers in a week because of the system restrictions of 4S + 3M ≤ 40 and 2S + 6M ≤ 50. The company will not be able to make these numbers in a week because of the system restrictions of 2S + 3S ≤ 40 and 4M + 6M ≤ 50. @superhelp101

OpenStudy (anonymous):

i think A

OpenStudy (anonymous):

or c

OpenStudy (superhelp101):

@A.K.C.15 i would choose A

OpenStudy (anonymous):

i was leaning towards it :) you're literally my fav on this site @superhelp101

OpenStudy (anonymous):

A system of equations is shown below: 2x = - y + 6 -4x + 3y = 8 What is the solution to this system of equations? (-1, -4) (1, 4) (4, 1) (-4, -1)

OpenStudy (superhelp101):

awhhhh OMG thank you for say that u amazing! :)

OpenStudy (superhelp101):

a really easy way to figuring this out is but plugging all of the coordinates in the equation and if X=X then the answer is right do u see what i mean??... if you don't let me know.

OpenStudy (anonymous):

i dont :(

OpenStudy (superhelp101):

okay so for example 2x = - y + 6 option A is (-1, -4) right u plug it in so it would look like 2(-1)= -(-4) +6 -2= 4+6 -2= 10 so A isn't right

OpenStudy (superhelp101):

are u getting it ?:)

OpenStudy (anonymous):

yes i thing c would be the answer

OpenStudy (superhelp101):

it is B 2x = - y + 6 -4x + 3y = 8 coordinates are (1, 4) 2(1)= -4 +6 2 = 2 YAS you should check each equation but i didn't lol so i say it is B

OpenStudy (anonymous):

Solve the following system: y = x + 3 4x + y = 18 6, 3) (3, 6) (-3, 6) (3, -6) i thing its B

OpenStudy (superhelp101):

whooo good! u getting this

OpenStudy (anonymous):

i have 7 more you don't have to finish helping if you don't want

OpenStudy (superhelp101):

it is 2;02 here haha but i will do my best :)

OpenStudy (superhelp101):

am

OpenStudy (anonymous):

haha ok thx

OpenStudy (anonymous):

Which pair of graphs should be plotted to show the solution to 5(1 over 2)x = 1 over 2(5)x? Graphs Q and R Graphs P and R Graphs R and S Graphs P and Q

OpenStudy (anonymous):

i think A

OpenStudy (superhelp101):

uh- huh yep!

OpenStudy (superhelp101):

okay so lets start out do you have an idea?

OpenStudy (anonymous):

legit no clue

OpenStudy (superhelp101):

:D

OpenStudy (anonymous):

huh?

OpenStudy (superhelp101):

so i kinda forgot about the graphs where u shade in stuff but i think that the answer might be A or C ;/

OpenStudy (anonymous):

which do you think more likely @superhelp101

OpenStudy (superhelp101):

;p i really don't want to get u this wrong but i am thinking A

OpenStudy (anonymous):

its ok

OpenStudy (anonymous):

Rex mixed x liters of fruit syrup with y liters of water to make a fruit punch. Three times the number of liters of water he mixed was 6 more than twice the number of liters of syrup he mixed. The total number of liters of syrup and water he mixed was 7 liters. Which graph best shows the number of liters of fruit syrup and water that Rex added to make the fruit punch?

OpenStudy (anonymous):

I do not see a graph, however, he added 3 liters of fruit syrup and 4 liters of water. so which would it be?

OpenStudy (anonymous):

i think D

OpenStudy (anonymous):

@superhelp101